### v1

# N2: A Unified Python Package and Test Bench for Nearest Neighbor-Based Matrix Completion

Abstract
Nearest neighbor (NN) methods have re-emerged as competitive tools for matrix completion, offering strong empirical performance and recent theoretical guarantees, including entry-wise error bounds, confidence intervals, and minimax optimality. Despite their simplicity, recent work has shown that NN approaches are robust to a range of missingness patterns and effective across diverse applications. This paper introduces N2, a unified Python package and testbed that consolidates a broad class of NN-based methods through a modular, extensible interface. Built for both researchers and practitioners, N2 supports rapid experimentation and benchmarking. Using this framework, we introduce a new NN variant that achieves state-of-the-art results in several settings. We also release a benchmark suite of real-world datasets—from healthcare and recommender systems to causal inference and LLM evaluation—designed to stress-test matrix completion methods beyond synthetic scenarios. Our experiments demonstrate that while classical methods excel on idealized data, NN-based techniques consistently outperform them in real-world settings.
1 Introduction
Nearest neighbor methods are a class of non-parametric algorithms widely used for regression, classification and pattern recognition. Due to their scalability and success under models with minimal assumptions, nearest neighbor methods have recently been adopted for practical fields such as matrix completion and counterfactual inference in panel data settings. Matrix completion is a well-established field that supplies practitioners with many tools to recover underlying matrices using partial or even noisy observations HMLZ (15); Cha (15); KMO (10), with recommendation systems KBV (09); Rec (11) as an important use-case. Panel data counterfactual inference aims at learning the treatment effect of policies across time Bai (09); BN (21); ABD+ (21). One important example is individualized healthcare predictions KSS+ (19). Nearest neighbor methods were recently recognized as effective in providing granular inference guarantees for both matrix completion and counterfactual inference when either the missingness or the policy treatment are not completely random and confounded MC (19); DTT+22a ; ADSS (23). They have also been recently leveraged to tackle distributional matrix completion settings FCAD (24); CFC+ (24)
Despite nearest neighbor methods popularity, there is no unified package that lets a user easily switch between different kinds of nearest neighbor algorithms for matrix completion and counterfactual inference. In this paper, we present a package ($\mathbf{N}^{2}$) to unify several nearest neighbor methods under a single interface, so users can easily choose the method that suits their data the best, for both scalar and distributional missing data problems. Next, we review two broad applications of nearest neighbors and matrix completion: recommendation systems and panel data settings.
Recommendation systems
In the problem of a recommendation system, multiple users submit ratings / preferences for a subset of items, and the vendor recommends items based on the partial rating information RS (05). Prediction of unseen ratings is the key to a good recommendation, and the predictions are based on the patterns of the partially observed ratings which can be formalized into a matrix completion problem. A popular instance is the Netflix problem, an open competition for the best prediction algorithm of user ratings of films. Users (rows of matrix) have the option to rate the films (columns of matrix). However, users typically rate only a subset of films, and the ratings are commonly biased towards their preferences. Under the common belief that only a few factors drive the preference and ratings of the users, many matrix completion algorithms have resorted to the low rank assumption on the ratings matrix Rec (11); CR (12). Variants of nearest neighbors (in short, $\mathsf{NN}$) recently were introduced to correct for the potential bias originating from the fact that observation patterns are influenced by the matrix entries (e.g. ratings) themselves and by unmeasured confounders ADSS (23); AADS (24).
Panel data
Panel data constitutes measurements of multiple units that are tracked over time (hence a type of longitudinal data) as they undergo different types of interventions at different times Bai (09). Panel data is used to analyze the causal effects of any policies, treatments, or business decisions that affect the subjects in the data, making this data type essential in econometrics and health studies. Matrix completion algorithms enable estimation of these causal effects at the unit×$\times$×time level, the most granular scale where causal effects can be estimated ABD+ (21); ADSS (23). For example, consider a mobile health app trying to learn the effectiveness of two exercise routines. Suppose the app alternates two routines across $N$ different users repeatedly over $T$ weeks, where their health activities (say physical step counts) throughout each week are recorded. Then, we can use matrix completion to impute the counterfactual step counts under treatment and control, allowing us to estimate causal effects at the most granular scale.
Our contributions
We make the following contributions: First, we present a unified framework for nearest neighbor algorithms that facilitates extending to new variants, especially for matrix completion problems. We leverage this framework to introduce then a new nearest neighbors algorithm, auto nearest neighbors (in short $\mathsf{AutoNN}$), which aims both to improve on the existing methods and demonstrate the ease of developing new variants with our library. Next, we introduce a unified, easy-to-implement nearest neighbor library that contains a breadth of nearest neighbor algorithms for matrix comple2tion problems. Next, we present a test bench called $\mathbf{N}^{2}$-Bench that comprises of several real-world data sets across applications. Finally, we illustrate our library’s wide applicability on the test benchmark.
1.1 Related work
We contextualize our contributions in the context of both nearest neighbors as a general algorithm and matrix completion specifically.
Nearest neighbors
As introduced above, nearest neighbor methods are widely used for non-parametric regression, classification, and pattern recognition CH (67); CBS (20). Recently, nearest neighbor methods were introduced as an effective algorithm for matrix completion problems LSSY (19); ADSS (23); DTT+22a , especially when the missingness depended on observations and unobserved confounders. The fact that nearest neighbor methods target a single entry at a time via matching makes them effective against various types of missing patterns. The class of algorithms has grown to account for a wide range of applications involving scalar or distributional settings; for instance, nearest neighbors are used for evaluating the personalized recommendation effect of a healthcare app on the average and the distribution of physical step counts DTT+22a ; CFC+ (24); FCAD (24); SPD (24) and are also used for policy evaluations ADSS (23). Our library, $\mathbf{N}^{2}$, is designed to easily implement the vanilla scalar versions of nearest neighbors LSSY (19); DTT+22a as well as its unweighted DTT+22b ; SPD (24) and weighted SPD (25) variants. Finally, $\mathbf{N}^{2}$ is also capable of distributional matrix completion CFC+ (24); FCAD (24).
Other matrix completion methods
Universal singular value thresholding (USVT), proposed in Cha (15); BC (22), is a classical spectral-based method for performing matrix completion; its core functionality is based on a singular value decomposition of the matrix and thresholding the singular values. SoftImpute, introduced by HMLZ (15), is another widely used optimization-based algorithm for matrix completion. The algorithm computes ridge-regression updates of the low-rank factors iteratively and finally soft-thresholds the singular values to impose a nuclear norm penalty. Notably USVT and SoftImpute have provable guarantees when missingness is completely at random, but empirically fail when the missing pattern depends on the observed entries or the unobserved confounders ADSS (23). Our real-world analysis in Sec. 4 once again demonstrates this point.
Existing software for matrix completion and nearest neighbors
Scikit-Learn PVG+ (11), a popular Python package for machine learning tools, implements a simple k-nearest neighbor algorithm for imputing missing values in a feature matrix. However, their implementation is designed for the feature matrix setting. So, neighbors are only defined across samples (row-wise). Additionally, they do not provide any implementation for more advanced nearest neighbor algorithms, nor does their package allow for easy extendability like our proposed package.
Organization
Sec. 2 contains an overview of the existing nearest neighbor algorithms implemented in our library, $\mathbf{N}^{2}$. Sec. 2.3 introduces a new nearest neighbor algorithm, $\mathsf{AutoNN}$, that is also implemented in $\mathbf{N}^{2}$. Sec. 3 touches upon the high-level (class) structure of $\mathbf{N}^{2}$, as well as the interface for practitioners. Sec. 4 tests different variants of nearest neighbor methods and classical methods on our new test bench of diverse datasets called $\mathbf{N}^{2}$-Bench. Finally, in Sec. 5, we provide concluding remarks and outline future directions of research.
2 Nearest Neighbors for Matrix Completion
We now introduce the mathematical model for matrix completion:
$\quad\text{for}\quad i\in[N],t\in[T]:\quad Z_{i,t}:=\begin{cases}X_{1}(i,t),...,X_{n}(i,t)\ \sim\ \mu_{i,t}&\quad\text{if}\quad A_{i,t}=1,\\ \mathrm{unknown}&\quad\text{if}\quad A_{i,t}=0.\end{cases}$ (1)
[…]
2.1 Unified framework
We introduce two general modules (namely Distance and Average) from which the variants of nearest neighbors are constructed. We introduce several shorthands used in the modules. Denote the collection of measurements, missingness, and weights:
$\mathcal{Z}:=\big{\{}Z_{j,s}\big{\}}_{j\in[N],s\in[T]},\quad\mathcal{A}:=\{A_{j,s}\}_{j\in[N],s\in[T]},\quad\text{and}\quad\mathcal{W}:=\{w_{j,s}\}_{j\in[N],s\in[T]}.$ (2)
Let $\varphi(x,x^{\prime})$ be a metric between $x,x^{\prime}\in\mathcal{X}$ for some space $\mathcal{X}$. Further define $\widehat{\varphi}(Z_{i,t},Z_{j,s})$ as a data-dependent distance between any two observed entries $(i,t)$ and $(j,s)$ of the matrix 1. The two modules can now be defined:
  • (i)
    $\textsc{Distance}(\widehat{\varphi},\mathcal{Z},\mathcal{A})$: Additional input is the data-dependent distance between entries of matrix $\widehat{\varphi}$ and output is the collection of row-wise and column-wise distance of matrix:
    $\rho_{i,j}^{\mathrm{row}}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}\widehat{\varphi}(Z_{i,s},Z_{j,s})}{\sum_{s\neq t}A_{i,s}A_{j,s}}\quad\text{and}\quad\rho_{t,s}^{\mathrm{col}}:=\frac{\sum_{j\neq i}A_{j,t}A_{j,s}\widehat{\varphi}(Z_{j,t},Z_{j,s})}{\sum_{j\neq i}A_{j,t}A_{j,s}},$ (3)
  • (ii)
    $\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$: Additional input are the weights $\mathcal{W}$, metric $\varphi$ and output is the optimizer
    $\widehat{\theta}=\mathop{\mathrm{missing}}{argmin}_{x\in\mathcal{X}}\sum_{j\in[N],s\in[T]}w_{j,s}A_{j,s}\varphi(x,Z_{j,s}).$ (4)
The Distance module calculates the row-wise and column-wise distance of the matrix, by taking the average of the observed entry-wise distance $\widehat{\varphi}(\cdot,\cdot)$. The Average module calculates the weighted average of observed measurements, where the notion of average depends on the metric $\varphi$ and the space $\mathcal{X}$ on which the metric $\varphi$ is defined. Notably, the weights $\mathcal{W}$ in the Average module encodes the entry information of the estimand.
Remark 1
[…]
2.2 Existing methods
We present existing variants of nearest neighbors using the two modules introduced Sec. 2.1; all the methods presented here are recovered by sequentially applying Distance and Average with the appropriate specification of $\widehat{\varphi}$, $\varphi$ and $\mathcal{W}$.
All methods except $\mathsf{AWNN}$ and our newly proposed $\mathsf{AutoNN}$, have binary weights i.e., $w_{j,s}\in\{0,1\}$. $\mathsf{AutoNN}$, detailed in Sec. 2.3, uses weights to carefully pool together the benefits of $\mathsf{TSNN}$ and $\mathsf{DRNN}$. $\mathsf{AWNN}$ SPD (25) improves upon $\mathsf{RowNN}$ by adaptively choosing the weights which optimally balances the bias-variance tradeoff of $\mathsf{RowNN}$ as follows
$\big{(}w_{1}^{\star}(i,t),...,w_{N}^{\star}(i,t)\big{)}:=\mathop{\mathrm{missing}}{argmin}_{(v_{1},...,v_{N})\in\Delta_{N}}2\log(2N)\widehat{\sigma}^{2}\sum_{k\in[N]}v_{k}^{2}+\sum_{k\in[N]}v_{k}A_{k,t}\rho_{i,k}^{\mathrm{row}}.$ (5)
where $\widehat{\sigma}^{2}$ is the estimated error and $\Delta_{N}$ is a simplex in N; see SPD (25) for details of 5. Tab. 1 contains a concise summary of the existing nearest neighbor variants; see App. B for a detailed exposition for each methods.
Table 1: Variants of nearest neighbors for matrix completion.
Type Method ${\widehat{\varphi}}(x,y)$
$\varphi(x,y)$
$w_{j,s}$
$n=1$ $\mathsf{RowNN}$  (LSSY, 19) (Alg. 1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{ColNN}$  (LSSY, 19) (Alg. 1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq 0,\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{TSNN}$  (SPD, 24) (Alg. 2) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{AWNN}$  (SPD, 25) (Alg. 5) $(x-y)^{2}$ $(x-y)^{2}$ $w_{j}^{\star}(i,t)\cdot\mathbf{1}({\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{DRNN}$  (DTT+22b, ) (Alg. 3) $\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$
$\mathsf{AutoNN}$  (Sec. 2.3) $\alpha\cdot\mathsf{DRNN}+(1-\alpha)\cdot\mathsf{TSNN}$
$n>1$ $\mathsf{KernelNN}$  (CFC+, 24) (Alg. 4) $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(x,y)$ $\mathsf{MMD}_{\mathbf{k}}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{W_{2}NN}$  (FCAD, 24) (Alg. 4) $\widehat{\mathsf{W}}_{2}^{2}(x,y)$ $\mathsf{W}_{2}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
Under the distributional matrix completion setting ($n>1$ in 1), the methods $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ in Tab. 1 take $\mu,\nu\in\mathcal{X}$ as square integrable probability measures, and $\varphi(\mu,\nu)$ as either the squared maximum mean discrepency (i.e. $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu,\nu)$, see MFS+ (17)) or squared Wasserstein metric (i.e., $\mathsf{W}_{2}(\mu,\nu)$, see Big (20)). Further, the entry-wise distance $\widehat{\varphi}(x,y)$ in this case is either the unbiased U-statistics estimator $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu_{i,t},\mu_{j,s})$ (see MFS+ (17)) or the quantile based estimator $\widehat{\mathsf{W}}_{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{W}_{2}(\mu_{i,t},\mu_{j,s})$ (see Big (20)).
2.3 New variant: Auto nearest neighbors
$\mathsf{TSNN}$ is a generalization of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ by setting one of the tuning parameters to zero (see Tab. 1), whereas the idea underlying $\mathsf{DRNN}$ is fundamentally different from that of $\mathsf{TSNN}$; $\mathsf{DRNN}$ debiases a naive combination of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ whereas $\mathsf{TSNN}$ simply boosts the number of measurements averaged upon, thereby gaining from lower variance. So we simply interpolate the two methods for some hyper-parameter $\alpha\in[0,1]$; see Tab. 1. Notably the hyper-parameter $\eta$ for both $\mathsf{DRNN}$ and $\mathsf{TSNN}$ are identical when interpolated.
Suppose $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ in 1 where $\varepsilon_{i,t}$ are centered i.i.d. sub-Gaussian distributions across $i$ and $t$. When $\sigma$ is large in magnitude, $\mathsf{TSNN}$ denoises the estimate by averaging over more samples, hence providing a superior performance compared to $\mathsf{DRNN}$ in a noisy scenario. When $\sigma$ is small so that bias of nearest neighbor is more prominent, $\mathsf{DRNN}$ effectively debiases the estimate so as to provide a superior performance compared to $\mathsf{TSNN}$. The linear interpolator $\mathsf{AutoNN}$ automatically adjusts to the underlying noise level and debiases or denoises accordingly; such property is critical when applying nearest neighbors to real world data set where the noise level is unknown. We refer to Fig. 1 for visual evidence.
Refer to caption Refer to caption
(a) High SNR (b) Low SNR
Figure 1: Error scaling for certain NN variants in synthetic experiments. See Sec. D.1 for details on the data-generating process and how the signal-to-noise ratio (SNR) is defined. Each point corresponds to the mean absolute error ± 1 standard error across 30 trials.
3 $\mathbf{N}^{2}$ Package and Interface
We now present our unified Python package, $\mathbf{N}^{2}$, for nearest neighbor algorithms for matrix completion. In particular, we provide a class structure which abstracts the estimation procedure utilized in each different nearest neighbor method as well as the the Distance and Average modules described above in Sec. 2. On top of that, our library facilitates easy extensions to other nearest neighbors algorithms and other data types on top of scalars and distributions. For example, as long as a distance and average notion are well defined, our library can be easily applied to a matrix of images or text strings.
Class structure.
The core functionality of $\mathbf{N}^{2}$ is based on two abstract classes: EstimationMethod and DataType.
EstimationMethod classes contain the logic to impute a missing entry such as how to use calculated distances. In the context of the modules in Sec. 2, this class determines the weighting function or, in the context of $\mathsf{DRNN}$ and $\mathsf{AutoNN}$, how to compose estimates. We separate this from the DataType abstraction because several estimation methods can be used for multiple data types. For example, RowRowEstimator implements the $\mathsf{RowNN}$ procedure for any data type given to it, such as scalars or distributions.
DataType classes implement the Distance and Average modules for any kind of data type (e.g. scalars which use squared distance and simple averaging). This abstract class allows for our package to extend to any data types beyond the ones we tested. For instance, a practitioner can easily add a DataType for text strings which uses vector embeddings to find distances and averages between between strings without needing to rewrite any of the estimation procedure.
Interface.
To use our library, a user simply has to instantiate a composite class NearestNeighborImputer with their EstimationMethod and DataType of choice. We provide constructor functions to automatically create popular NearestNeighborImputer classes such as a two-sided nearest neighbor estimator with the scalar data type. From a design pattern point of view, this is known as a Composite design pattern (GHJV, 93, pg. 163). We use this design pattern so that anyone looking to customize the estimation procedure can do so for any kind of data type simultaneously. Similarly, with the exception of doubly robust estimators, each estimation procedure works out of the box with any data type that implements the DataType abstract class. The Doubly robust estimation method does not work out of the box with distributions because a subtraction operation is not well defined in the distribution space.
Finally, the user simply needs to input (i) a data matrix, (ii) a mask matrix which specifies which values are missing, and (iii) the row and column to impute. Thus, a user can test out different estimation procedures by changing just one line of code. Separately from the core functionality, we have also implemented several cross-validation classes which take in a NearestNeighborImputer class and find the best hyperparameters to use (e.g., distance thresholds and weights).
4 $\mathbf{N}^{2}$-Bench and Results
In this section, we evaluate several nearest neighbor algorithms provided by our library, $\mathbf{N}^{2}$, on real-world data. As part of our package, we include data loaders which automatically download the necessary datasets and format them for evaluation. These datasets and loaders comprise our proposed benchmark for nearest neighbor matrix completion algorithms, $\mathbf{N}^{2}$-Bench. We also test several existing popular matrix completion techniques (HMLZ (15); Cha (15)). For details on our experimental setup, computing hardware, and boxplot generation, see App. D.
4.1 Personalized healthcare: HeartSteps
The HeartSteps V1 study (HeartSteps study for short) is a clinical trial designed to measure the efficacy of the HeartSteps mobile application for encouraging non-sedentary activity KSS+ (19). The HeartSteps V1 data and its subsequent extensions have been widely used for benchmarking a variety of tasks including counterfactual inference of treatment effect DTT+22a ; CFC+ (24), reinforcement learning for intervention selection LGKM (20), and micro-randomized trial design QWC+ (22). In the HeartSteps study, $N=37$ participants were under a 6-week period micro-randomized trial, where they were provided with a mobile application and an activity tracker. Participants independently received a notification with probability $p=0.6$ for $5$ pre-determined decision points per day for 40 days ($T=200$). We denote observed entries $Z_{i,t}$ as the mean participant step count for one hour after a notification was sent and unobserved entries as the unknown step count for decision points where no notification was sent. Our task is to estimate the counterfactual outcomes: the participant’s step count should they have received a different treatment (notification or no notification) than they did at specific time points during the study.
Results & Discussion.
We benchmark the performance of the matrix completion methods by measuring absolute error on held-out observed step counts across 10 participants in the last 50 decision points. We use the remaining data to find nearest neighbor hyperparameters using cross-validation. To benchmark the distributional nearest neighbors methods ($\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$) against the scalar methods, we first set each entry to have the number of samples $n=60$, where each sample is the 1 minute step count before imputation. Then, we take the mean of the imputed empirical distribution as the estimate.
In Fig. 2(a), we compare the absolute error of the imputed values across the nearest neighbor and baseline methods. The scalar nearest neighbor methods far out-perform USVT and are on par with SoftImpute. The two distributional nearest neighbor methods far outperform all methods operating in the scalar setting; it suggests that matching by distributions collect more homogeneous neighbors, thereby decreases the bias of the method, compared to matching only the first moments as done in most scalar matrix nearest neighbor methods.
In Fig. 2 panel (b), we show an example of an imputed entry in the distributional nearest neighbors setting. In this case, the ground truth distribution is bimodal, as the participant was largely sedentary (0 steps) with small amounts of activity. While both $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ capture the sedentary behavior of the participant, $\mathsf{KernelNN}$ is able to recover the bimodality of the original distribution whereas $\mathsf{W_{2}NN}$ cannot.
Refer to caption Refer to caption
(a) Absolute error of mean step count prediction (b) $\mathsf{KernelNN}$ vs. $\mathsf{W_{2}NN}$
Figure 2: HeartSteps: estimating step count under scalar and distributional matrix completion settings. Panel (a) shows the absolute error of predicted step count of the nearest neighbor methods against matrix completion baselines (SoftImpute, USVT). Panel (b) shows an example of an imputed entry in the distributional matrix completion setting.
4.2 Movie recommendations: MovieLens
Figure 3: MovieLens: Estimation error for a random subsample of size 500. For experimental settings and discussion see Sec. 4.2.
The MovieLens 1M dataset HK (15) contains 1 million ratings (1–5 stars) from 6,040 users on 3,952 movies. Collaborative filtering on MovieLens has long been a benchmark for matrix-completion methods: neighborhood-based algorithms SKKR (01), latent-factor models KBV (09), and, more recently, nearest neighbors interpreted as blind regression under a latent–variable model LSSY (19). These assist practitioners in data-driven recommendation systems, since more accurate rating imputation directly drives better personalized suggestions and user engagement. This is a standard scalar matrix completion problem with $N=6,040$ and $T=3,952$. Each rating is an integer in $\{1,\dots,5\}$. The dataset has a very high percentage of missing values: 95.53% missing. Our task is to estimate unobserved ratings using various matrix completion algorithms. We benchmark the performance of nearest neighbors against matrix factorization by measuring absolute error on held-out ratings. See Sec. D.3 for additional details on the dataset.
Results & Discussion.
We fit the nearest neighbor methods using a random sample of size 100 from the first 80% of the dataset to choose nearest neighbor hyperparameters via cross-validation.We then test the method on a random subsample of size 500 from the last 20% of the dataset. As observed in Fig. 3, all nearest neighbor methods have a lower average error than USVT and a much lower standard deviation of errors, with $\mathsf{ColNN}$, $\mathsf{RowNN}$, $\mathsf{DRNN}$, and $\mathsf{AutoNN}$ performing the best out of the nearest neighbor methods. SoftImpute performs on par with the nearest neighbor methods. Note that the nearest neighbor methods perform well even while only being trained on a tiny subset of the data of size 100 out of the 1 million ratings available.
4.3 Counterfactual inference for panel data: Proposition 99
Refer to caption Refer to caption
(a) Synthetic controls for California (b) Absolute error on control states
in post-intervention period
Figure 4: Nearest neighbor methods generate high-fidelity synthetic controls in counterfactual inference for panel data. For exact settings and further discussion see Sec. 4.3.
Next we consider a panel data setting, where our goal is to estimate the effect of the California Tobacco Tax and Health Protection Act of 1988 (a.k.a. Proposition 99) on annual state-level cigarette consumption111measured as per capita cigarette sales in packs. By definition, the counterfactual cigarette consumption in California—had Proposition 99 never been enacted—is not observed. ADH (10) introduce the notion of a “synthetic control” to serve as a proxy for this unobserved value based on “neighboring” control states that never instituted a tobacco tax. These states are not close in a geographical sense, but rather close due to similarities in other covariates [^2] GDP per capita, beer consumption, percent aged 15–24, and cigarette retail prices. We take a different approach and use only the observed cigarette consumption levels from the control states, of which there are 38 in total. Thus, we frame our problem as a scalar matrix completion problem with $N=39$ and $T=31$ (see 1). The last row in the matrix corresponds to the state of California.
Results & Discussion.
For each method, we use a 64-16-20 train-validation-test split and use cross validation to fit any hyperparameters. Fig. 4 plots the various synthetic controls for California (left) and absolute error of each method on the 38 control states, for which we do observe the no-treatment values (right). From Fig. 4(a), we see that nearest neighbor methods, in particular $\mathsf{TSNN}$ and $\mathsf{RowNN}$, are roughly on par with the gold-standard synthetic control method of ADH (10) (“SC”) for estimating California’s counterfactual cigarette consumption in the post-intervention period (after 1989). This is despite the fact that the nearest neighbor methods rely on less information for the estimation task. From Fig. 4(b), we see that all nearest neighbor methods, with the exception of $\mathsf{ColNN}$, achieve similar error levels as the synthetic control baseline. $\mathsf{RowNN}$ achieves even lower error levels. See supplementary experiment details in Sec. D.4.
4.4 Efficient LLM evaluation: PromptEval
The rapid advancement of LLMs have placed them at the center of many modern machine learning systems, from chatbots to aids in medical education GHC+ (25). In practice, system architects want to strike the right balance of real-world performance and cost, but navigating this Pareto frontier is a daunting task. 2024 alone saw at least 10 new models from Anthropic, Google, Meta, and OpenAI, not even counting the multitude of open-source fine-tuned models built on top of these. On specific tasks, smaller, fine-tuned models may even outperform the latest frontier models, in addition to being more cost effective.
We investigate how matrix completion, specifically nearest neighbor methods, can alleviate some of these burdens. We use the PromptEval dataset PXW+ (24), which evaluates $15$ open-source language models (ranging in size from 3B to 70B parameters) and $100$ different prompting techniques across the 57 tasks of the MMLU benchmark (HBB+, 20). In practice, the performance of a model depends—sometimes dramatically—on the precise input prompt. This suggests that we need to consider the performance of a model across a wide range of prompts, rather than any one prompt in particular. Thus, we model this problem as a distributional matrix completion problem with $N=15$, $T=57$, and $n=100$. Given one of 57 tasks, we aim to accurately characterize the performance of each model without resorting to exhaustive evaluation. Nearest neighbors leverage commonalities across models and tasks to estimate the performance distribution of each entry, which was otherwise not considered in PXW+ (24); previous literature achieves efficient evaluation per model and task in isolation without leveraging any across model / task information.
Refer to caption Refer to caption
(a) Mean KS distance between estimated (b) $\mathsf{KernelNN}$ vs. $\mathsf{W_{2}NN}$
and ground-truth distributions
Figure 5: Distributional nearest neighbor methods enable efficient LLM evaluation on MMLU. We estimate LLM score distributions across all models and tasks given only a limited number of model-task evaluations, determined by the propensity $p$. See Sec. 4.4 for a detailed discussion.
Results & Discussion.
We randomly include each entry in the matrix independently with probability $p\in\mathopen{}\mathclose{{}\left\{0.3,0.5,0.7}\right\}$ and impute the missing entries using the $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ methods of Tab. 1. For each method, we consider both the the row-wise and column-wise variants. Fig. 5(a) reports the mean Kolmogorov-Smirnov (KS) distance between the imputed and ground-truth distributions across the entries in the test set for varying missingness values. As expected, estimation error decreases as $p$ increases. Fig. 5(b) visualizes the imputed distributions using row-wise $\mathsf{KernelNN}$ and column-wise $\mathsf{W_{2}NN}$ (at $p=0.7$) for a select entry, along with the ground-truth distribution. Even with 30% of matrix entries missing, distributional NN methods are able to recover the underlying distribution.
5 Conclusion
In this paper, we present a unified framework, Python library ($\mathbf{N}^{2}$), and test bench ($\mathbf{N}^{2}$-Bench) for nearest neighbor-based matrix completion algorithms for both scalar and distributional settings. We demonstrate how our library supports a diverse set of datasets spanning recommendation systems (MovieLens), patient-level healthcare causal inference (HeartSteps), counterfactual inference for panel data (Proposition 99), and LLM evaluation (PromptEval). Our framework and library facilitate researchers and practitioners to easily try out different nearest neighbor methods on their dataset of choice as well as extend the library to more complex nearest neighbor methods.
Several future directions are natural from our work. Here we focused on unweighted and specific weighing schemes; several other weighting strategies can be easily incorporated into the library. Next, being able to deal with larger matrices requires handling distributed datasets as well as speeding up the runtime of $\mathbf{N}^{2}$, e.g., via efficient implementations using matrix multiplication or approximate nearest neighbors strategies.
Appendix A Structural assumptions
[…]
The first assumption specifies the factor structure on the mean; that is, there exists latent factors $u_{i},v_{t}$ that collectively characterize the signal of each entry $(i,t)$ of the matrix [37, 20, 4, 16, 23]. Such a factor model is analogous to the low rank assumptions commonly imposed in matrix completion [19]. The second assumption specifies how the missing pattern $A_{i,t}$ was generated; for instance missing completely at random (MCAR) assumes that $A_{i,t}$ are independent to all other randomness present in the model 1 and that all entries have positive probability of being observed.
A.1 Factor model
For the scalar matrix completion problem, i.e., 1 with $n=1$, the main goal is to learn (or impute) the mean of the underlying distribution $\theta_{i,t}$ for any missing entries [37, 21, 20, 4, 48, 49]. The majority of this literature assumes (i) an additive noise model $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ for centered i.i.d. sub-Gaussian noise $\varepsilon$ and (ii) mean factor model, i.e., $\theta_{i,t}=f(u_{i},v_{t})$ for some latent factors $u_{i},v_{t}$ and real valued function $f$.
[…]
A.2 Missingness pattern
For both the scalar and distributional matrix completion problem 1, the missing pattern (i.e., how the missingness $A_{j,s}$ was generated) can be categorized into three classes using the taxonomy of [36]: missing-completely-at-random (MCAR), missing-at-random (MAR) and missing-not-at-random (MNAR). MCAR assumes that the missingness $A_{i,t}$ is exogenous (independently generated from all the randomness in the model) and i.i.d. with propensity ${\mathbb{P}}(A_{i,t}=1)=p>0$ for all $(i,t)$. MAR is a more challenging scenario compared to MCAR as missingness is not exogenous, but its randomness depends on the observations. Further, propensities $p_{i,t}$ may differ for entries $(i,t)$ but positivity still holds, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}>0$. An important instance for MAR is the adaptive randomized policies [22]. The MNAR setup is the most challenging as it assumes the missingness depends on the unobserved latent confounders, while positivity may also be violated, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}=0$. The staggered adoption pattern, where a unit remains treated once a unit is treated at some adoption time, is a popular example of MNAR, mainly because positivity is violated. See [2, 5] for more details on staggered adoption.
We briefly outline the structural assumptions existing nearest neighbor methods were shown to work with provable guarantees; for all the existing methods, factor models (with slightly different details; compare the mean factorization [37] and the distribution factorization [16, 23]) are all commonly assumed.
  • (Scalar matrix completion) The vanilla versions of nearest neighbors ($\mathsf{RowNN}$) in [37, 20] are shown to work for MCAR and MAR setup; the latter shows that simple nearest neighbors can provably impute the mean when the missingness is fully adaptive across all users and history. The variants of vanilla nearest neighbors $\mathsf{DRNN}$ [21] is proven to work under MCAR, while $\mathsf{TSNN}$ [48] is proven to work under unobserved confounding, i.e., MNAR.
  • (Distributional matrix completion) The $\mathsf{KernelNN}$ [16] is shown to recover the underlying distribution under MNAR, whereas $\mathsf{W_{2}NN}$ [23] is shown to work under MCAR.
Appendix B Nearest neighbor algorithms
The nearest neighbor methods introduced in Tab. 1 are elaborated in this section. We present two versions of each method; the first version explicitly constructs neighborhoods instead of subtly embedding them in the weights $\mathcal{W}$ of the Average module, and the second version specifies how each methods can be recovered by applying the two modules, Distance and Average, sequentially.
B.1 Vanilla nearest neighbors
We elaborate on the discussion in Rem. 1 and provide here a detailed algorithm based on the explicit construction of neighborhoods, which is essentially equivalent to $\mathsf{RowNN}$ in Tab. 1. The inputs are measurements $\mathcal{Z}$, missingness $\mathcal{A}$, the target index $(i,t)$, and the radius $\eta$.
  • Step 1:
    (Distance between rows) Calculate the distance between row $i$ and any row $j\in[N]\setminus\{i\}$ by averaging the squared Euclidean distance across overlapping columns:
    $\rho_{i,j}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}(Z_{i,s}-Z_{j,s})^{2}}{\sum_{s\neq t}A_{i,s}A_{j,s}}.$ (6)
  • Step 2:
    (Construct neighborhood) Construct a neighborhood of radius $\eta$ within the $t$th column using the distances $\{\rho_{i,j}:j\neq i\}$:
    $\mathbf{N}_{t,\eta}:=\big{\{}j\in[N]\setminus\{i\}:\rho_{i,j}\leq\eta\big{\}}$ (7)
  • Step 3:
    (Average across observed neighbors) Take the average of measurements within the neighborhood:
    $\widehat{\theta}_{i,t,\eta}:=\frac{1}{|\mathbf{N}_{t,\eta}|}\sum_{j\in\mathbf{N}_{t,\eta}}A_{j,t}Z_{j,t}.$ (8)
In practice, the input $\eta$ for $\mathsf{RowNN}$ should be optimized via cross-validation; we refer the reader to App. C for a detailed implementation.
We specify the exact implementation of the two modules Distance, Average to recover $\mathsf{RowNN}$:
Input: $\mathcal{Z},\mathcal{A},\eta,$ $(i,t)$
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
2 Initialize hyper-parameter ${\eta}\leftarrow(\eta_{1},0)$
3 Calculate row-wise metric $\big{\{}\rho_{i,j}^{\mathrm{row}}:j\neq i\big{\}}\leftarrow$ $\textsc{Distance}(\widehat{\varphi},\mathcal{Z},\mathcal{A})$
4 Initialize weight $w_{j,s}\leftarrow\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$ return $\widehat{\theta}_{i,t}$
Algorithm 1 $\mathsf{RowNN}$ for scalar nearest neighbor
The discussion for $\mathsf{RowNN}$ here can be identically made for $\mathsf{ColNN}$ as well.
B.2 Two-sided and doubly-robust nearest neighbors
We elaborate on the variants of the vanilla nearest neighbors algorithm $\mathsf{TSNN}$ and $\mathsf{DRNN}$ in Tab. 1; we first elaborate on an equivalent version of each of the methods which explicitly constructs neighborhoods.
In the following three step procedure, $\mathsf{DRNN}$ and $\mathsf{TSNN}$ differs in the last averaging step: the inputs are the measurements $\mathcal{Z}$, missingness $\mathcal{A}$, the target index $(i,t)$, and the radii $\eta=(\eta_{1},\eta_{2})$.
  • Step 1:
    (Distance between rows) Calculate the distance between row $i$ and any row $j\in[N]\setminus\{i\}$ and the distance between column $t$ and any column $s\in[T]\setminus\{t\}$:
    $\rho_{i,j}^{\mathrm{row}}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}(Z_{i,s}-Z_{j,s})^{2}}{\sum_{s\neq t}A_{i,s}A_{j,s}}\quad\text{and}\quad\rho_{t,s}^{\mathrm{col}}:=\frac{\sum_{j\neq i}A_{j,t}A_{j,s}(Z_{j,t}-Z_{j,s})^{2}}{\sum_{j\neq i}A_{j,t}A_{j,s}}$ (9)
  • Step 2:
    (Construct neighborhood) Construct a row-wise and column-wise neighborhood of radius $\eta_{1}$ and $\eta_{2}$ respectively,
    $\mathbf{N}_{t,\eta_{1}}^{\mathrm{row}}:=\big{\{}j\in[N]\setminus\{i\}:\rho_{i,j}^{\mathrm{row}}\leq\eta\big{\}}\quad\text{and}\quad\mathbf{N}_{i,\eta_{2}}^{\mathrm{col}}:=\big{\{}s\in[T]\setminus\{t\}:\rho_{t,s}^{\mathrm{col}}\leq\eta\big{\}}$ (10)
  • Step 3:
    (Average across observed neighbors) Take the average of measurements within the neighborhood; the first and the second averaging correspond to $\mathsf{DRNN}$ and $\mathsf{TSNN}$ respectively:
    $\widehat{\theta}_{i,t,\eta}^{\mathsf{DR}}:=\frac{\sum_{j\in\mathbf{N}_{t,\eta_{1}}^{\mathrm{row}},s\in\mathbf{N}_{i,\eta_{2}}^{\mathrm{col}}}A_{j,t}A_{i,s}A_{j,s}\big{(}Z_{j,t}+Z_{i,s}-Z_{j,s}\big{)}}{\sum_{j\in\mathbf{N}_{t,\eta_{1}}^{\mathrm{row}},s\in\mathbf{N}_{i,\eta_{2}}^{\mathrm{col}}}A_{j,t}A_{i,s}A_{j,s}}\quad\text{and}\quad$ (11)
    $\widehat{\theta}_{i,t,\eta}^{\mathsf{TS}}:=\frac{\sum_{j\in\mathbf{N}_{t,\eta_{1}}^{\mathrm{row}},s\in\mathbf{N}_{i,\eta_{2}}^{\mathrm{col}}}A_{j,s}Z_{j,s}}{\sum_{j\in\mathbf{N}_{t,\eta_{1}}^{\mathrm{row}},s\in\mathbf{N}_{i,\eta_{2}}^{\mathrm{col}}}A_{j,s}}.$ (12)
Next, we specity the exact implemention of the two modules Distance and Average to recover $\mathsf{TSNN}$ and $\mathsf{DRNN}$:
Input: $\mathcal{Z},\mathcal{A},\eta,$ $(i,t)$
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
2 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},\eta_{2})$
3 Calculate row-wise and column-wise metric $\big{\{}\rho_{i,j}^{\mathrm{row}}:j\neq i\big{\}},\big{\{}\rho_{t,s}^{\mathrm{col}}:s\neq t\big{\}}\leftarrow$ $\textsc{Distance}(\widehat{\varphi},Z,A)$
4 Initialize weight $w_{j,s}\leftarrow\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$ return $\widehat{\theta}_{i,t}$
Algorithm 2 $\mathsf{TSNN}$ for scalar matrix completion
For $\mathsf{DRNN}$ algorithm below, we consider $\mathcal{Z}$ and $\mathcal{A}$ to be $N\times T$ sized matrices, so that their transpose is well defined. Then note that $\mathsf{ColNN}$ is simply applying Alg. 1 with transposed observation matrices.
Input: $\mathcal{Z},\mathcal{A},\eta,$ $(i,t)$
1 Initialize $\mathsf{RowNN}\leftarrow$ Alg. 1 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
2 Initialize $\mathsf{ColNN}\leftarrow$ Alg. 1 with input $(\mathcal{Z}^{T},\mathcal{A}^{T},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
3 Initialize $\mathsf{TSNN}\leftarrow$ Alg. 2 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and ${\eta}\leftarrow(\eta_{1},\eta_{2})$
Calculate $\widehat{\theta}_{i,t}\leftarrow\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$ return $\widehat{\theta}_{i,t}$
Algorithm 3 $\mathsf{DRNN}$ for scalar matrix completion
B.3 Distributional nearest neighbors
Unlike the scalar nearest neighbor methods, distributional nearest neighbors necessitate a distributional notion of distance between rows and columns of matrix and a distributional analog of averaging. [16] and [23] use maximum mean discrepency (in short $\mathsf{MMD}$) of kernel mean embeddings [40] and Wasserstein metric (in short $\mathsf{W}_{2}$[10] respectively both for defining the distance between rows / columns and for averaging. The corresponding barycenters of $\mathsf{MMD}$ and $\mathsf{W}_{2}$ [14, 9] are used for averaging, and so the methods are coined kernel nearest neighbors (in short $\mathsf{KernelNN}$) and Wasserstein nearest neighbors (in short $\mathsf{W_{2}NN}$) respectively.
We elaborate on a vanilla version three step procedure of $\mathsf{KernelNN},\mathsf{W_{2}NN}$ that explicitly constructs neighborhoods. The input are measurements $\mathcal{Z}$, missingness $\mathcal{A}$, the target index $(i,t)$ and the radius $\eta$,
  • Step 1:
    (Distance between rows) Calculate the distance between row $i$ and any row $j\in[N]\setminus\{i\}$ by averaging the estimator of distribution metric $\widehat{\varrho}$:
    $\rho_{i,j}^{\mathsf{MMD}}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(Z_{i,s},Z_{j,s})}{\sum_{s\neq t}A_{i,s}A_{j,s}}\quad\text{and}\quad\rho_{i,j}^{\mathsf{W}_{2}}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}\widehat{\mathsf{W}}_{2}^{2}(Z_{i,s},Z_{j,s})}{\sum_{s\neq t}A_{i,s}A_{j,s}}.$ (13)
  • Step 2:
    (Construct neighborhood) Construct a neighborhood of radius $\eta$ within the $t$th column using the distances $\{\rho_{i,j}:j\neq i\}$:
    $\mathbf{N}_{t,\eta}^{\mathsf{MMD}}:=\big{\{}j\in[N]\setminus\{i\}:\rho_{i,j}^{\mathsf{MMD}}\leq\eta\big{\}}\quad\text{and}\quad\mathbf{N}_{t,\eta}^{\mathsf{W}_{2}}:=\big{\{}j\in[N]\setminus\{i\}:\rho_{i,j}^{\mathsf{W}_{2}}\leq\eta\big{\}}$ (14)
  • Step 3:
    (Average across observed neighbors) Set $\mu^{Z}_{i,t}=n^{-1}\sum_{\ell=1}^{n}\boldsymbol{\delta}_{X_{\ell}(i,t)}$ as the empirical measure of the multiple measurements $Z_{i,t}$. Take the barycenter within the neighborhood:
    $\widehat{\mu}^{\mathsf{MMD}}_{i,t,\eta}$ $:=\frac{1}{|\mathbf{N}_{t,\eta}^{\mathsf{MMD}}|}\sum_{j\in\mathbf{N}_{t,\eta}^{\mathsf{MMD}}}A_{j,t}\mu_{j,t}^{Z}\quad\text{and}$ (15)
    $\widehat{\mu}^{\mathsf{W}_{2}}_{i,t,\eta}$ $:=\mathop{\mathrm{missing}}{argmin}_{\mu}\sum_{j\in\in\mathbf{N}_{t,\eta}^{\mathsf{W}_{2}}}\mathsf{W}_{2}^{2}(\mu,\mu_{j,t}^{Z}).$ (16)
For further details on the $\mathsf{W}_{2}$ and $\mathsf{MMD}$ algorithms see [23] and [16], respectively.
Input: $\mathcal{Z},\mathcal{A},\mathbf{k},\eta,$ $(i,t)$
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow$ $\widehat{\mathsf{MMD}}^{2}_{\mathbf{k}}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$ or $\widehat{\mathsf{W}}_{2}^{2}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$
2 Initialize metric $\varphi(x,y)\leftarrow$ ${\mathsf{MMD}}^{2}_{\mathbf{k}}(x,y)$ or $\mathsf{W}_{2}^{2}(x,y)$
3 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},0)$
4 Calculate row-wise metric $\big{\{}\rho_{i,j}^{\mathrm{row}}:j\neq i\big{\}}\leftarrow$ $\textsc{Distance}(\widehat{\varphi},Z,A)$
5 Initialize weight $w_{j,s}\leftarrow\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
Calculate average $\widehat{\mu}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$ return $\widehat{\mu}_{i,t}$
Algorithm 4 Vanilla (row-wise) distributional nearest neighbor
B.4 Adaptively weighted nearest neighbors
We elaborate on the adaptive variant of the vanilla nearest neighbor algorithm $\mathsf{AWNN}$ as mentioned in Sec. 2.2 and Tab. 1. The input are measurements $\mathcal{Z}$, and missingness $\mathcal{A}$. Note that there is no need for radius parameter $\eta$ and hence no CV.
  • Step 1:
    (Distance between rows and initial noise variance estimate) Calculate an estimate for noise variance and then the distance between any pair of distinct rows $i,j\in[N]$ by averaging the squared Euclidean distance across overlapping columns:
    $\rho_{i,j}:=\frac{\sum_{s\neq t}A_{i,s}A_{j,s}(Z_{i,s}-Z_{j,s})^{2}}{\sum_{s\neq t}A_{i,s}A_{j,s}},\qquad\overline{Z}\leftarrow\frac{\sum_{j,s\in[N]\times[T]}A_{j,s}Z_{j,s}}{\sum_{j,s\in[N]\times[T]}A_{j,s}},$ (17)
    $\widehat{\sigma}^{2}\leftarrow\frac{\sum_{j,s\in[N]\times[T]}A_{j,s}(Z_{j,s}-\overline{Z})^{2}}{\sum_{j,s\in[N]\times[T]}A_{j,s}}$ (18)
  • Step 2:
    (Construct weights) For all rows and columns $(i,t)\in[N]\times[T]$, evaluate $w^{(i,t)}=(w_{1,t},\cdots,w_{n,t})$, the weights that optimally minimizes the following loss involving an estimate of the noise variance $\widehat{\sigma}^{2}$,
    $w^{(i,t)}=\mbox{arg min}_{\widehat{w}^{(i,t)}}\mathopen{}\mathclose{{}\left[2\log(2m/\delta)\widehat{\sigma}^{2}\|\widehat{w}^{(i,t)}\|_{2}^{2}+\sum_{i^{\prime}\in[N]}\widehat{w}_{i^{\prime},t}A_{i^{\prime},t}\widehat{\rho}_{i^{\prime},i}}\right],$ (19)
    where $\widehat{w}^{(i,t)}=(\widehat{w}_{1,t},\cdots,\widehat{w}_{n,t})$ is a non-negative vector that satisfy $\sum_{i^{\prime}=1}^{n}\widehat{w}_{i^{\prime},t}A_{i^{\prime},t}=1$.
  • Step 3:
    (Weighted average) Take the weighted average of measurements:
    $\widehat{\theta}_{i,t}=\sum_{i^{\prime}\in[N]}\widehat{w}_{i^{\prime},t}A_{i^{\prime},t}X_{i^{\prime},t},\qquad\forall(i,t)\in[N]\times[T]$ (20)
  • Step 4:
    (Fixed point iteration over noise variance) Obtain new estimate of noise variance and stop if difference between old and new $\widehat{\sigma}^{2}$ is small.
    $\widehat{\sigma}^{2}\leftarrow\frac{1}{\sum_{i\in[N],t\in[T]}A_{i,t}}\sum_{i\in[N],t\in[T]}\mathopen{}\mathclose{{}\left(Z_{i,t}-\widehat{\theta}_{i,t}}\right)^{2}A_{i,t}$ (21)
No cross-validation in AWNN
The optimization problem in 19 can be solved exactly in linear time (worst case complexity) using convex optimization [49]. $\mathsf{AWNN}$ doesn’t rely on radius parameter $\eta$. Not only it automatically assigns neighbors to ${(i,t)^{th}}$ entry during its weight calculation(non-neighbors get zero weight), but also takes into account the distance of the neighbors from the ${(i,t)^{th}}$ entry. The closer neighors get higher weights and vice - versa.
We further specify the exact implementation of the two modules Distance, Average to recover $\mathsf{AWNN}$:
Input: $\mathcal{Z},\mathcal{A},$ $(i,t)$
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
2 Initialize noise - variance estimate ${\sigma_{\epsilon}^{2}}\leftarrow\mathrm{Variance}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\{Z_{i,t}}\right\}_{(i,t)\in[N]\times[T]}}\right)$
3 Calculate row-wise metric $\big{\{}\rho_{i,j}^{\mathrm{row}}:j\neq i\big{\}}\leftarrow$ $\textsc{Distance}(\widehat{\varphi},\mathcal{Z},\mathcal{A})$
4 Initialize weight $\mathopen{}\mathclose{{}\left\{w_{1,t},\dots,w_{n,t}}\right\}\leftarrow\mbox{arg min}_{\widehat{w}^{(i,t)}}\mathopen{}\mathclose{{}\left[2\log(2m/\delta)\widehat{\sigma}^{2}\|\widehat{w}^{(i,t)}\|_{2}^{2}+\sum_{i^{\prime}\in[N]}\widehat{w}_{i^{\prime},t}A_{i^{\prime},t}\widehat{\rho}_{i^{\prime},i}}\right]$
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$ return $\widehat{\theta}_{i,t}$
Algorithm 5 $\mathsf{AWNN}$ for scalar nearest neighbor
Appendix C Cross-Validation
For each nearest neighbor method, we use cross-validation to optimize hyperparameters including distance thresholds and weights, depending on which nearest neighbor algorithm is chosen. Specifically, for each experiment, we choose a subset of the training test to optimize hyperparameters by masking those matrix cells and then estimating the masked values. We utilize the HyperOpt library [13] to optimize (possibly multiple) hyperparamters using the Tree of Parzen Estimator [7], a Bayesian optimization method. Our package supports both regular distance thresholds and percentile-based thresholds, which adapt to the distances calculated within the specific dataset.
Appendix D Case Study Details
The boxplots are generated using matplotlib’s [30] standard boxplot function. The box shows the first, second, and third quartiles. The bottom line shows the first quartile minus the 1.5$\times$ the interquartile range. The top line shows the third quartile plus 1.5$\times$ the interquartile range. All experiments are run on standard computing hardware (MacBook Pro with an M2 Pro CPU with 32 GB of RAM).
D.1 Synthetic data generation
Generate $Z_{i,t}=X_{i,t}\sim N(\theta_{i,t},\sigma^{2})$, i.e., scalar matrix completion setting, with a linear factor structure $\theta_{i,t}=u_{i}v_{t}$. Row latent factors $u_{i}\in^{4}$ are i.i.d. generated across $i=1,...,N$, where each entry of $u_{i}$ follow a uniform distribution with support $[-0.5,0.5]$; column latent factors $v_{t}\in^{4}$ are generated in an identical manner. The missingness is MCAR with propensity $p_{i,t}=0.5$ for all $i$ and $t$. Further, the size of column and rows are identical $N=T$. For the left panel in Fig. 1, the noise level is set as $\sigma=0.001$ and for the right panel $\sigma=1$.
D.2 HeartSteps V1
The mobile application was designed to send notifications to users at various times during the day to encourage anti-sedentary activity such as stretching or walking. Participants could be marked as unavailable during decision points if they were in transit or snoozed their notifications, so notifications were only sent randomly if a participant was available and were never sent if they were unavailable. To process the data in the framework of 1, we let matrix entry $Z_{i,t}$ be the average one hour step count for participant $i$ and decision point $t$ when a notification is sent (i.e. $A_{i,t}=1$) and unknown when a notification is not sent (i.e. $A_{i,t}=0$). The treatment assignment pattern is represented as the 37 x 200 matrix visualized in Fig. 6. We use the dataset downloaded from https://github.com/klasnja/HeartStepsV1 (CC-BY-4.0 License).
Figure 6: HeartSteps V1 data notification pattern. The dark blue entries indicate that the app sent a notification to a sedentary participant—the entry has value $A_{i,t}=1$. The white entries indicate that the participant was available but did not receive a notification or they were active immediately prior to the decision point. The light blue entries indicate the participant was unavailable. We assign the value $A_{i,t}=0$ for all the white and light blue entries.
D.3 MovieLens
We load MovieLens via a custom MovieLensDataLoader that (i) downloads and caches the ml-1m.zip archive, (ii) reads ratings.dat into a user $\times$ movie pivot table, and (iii) constructs the binary mask where observed entries correspond to rated user–movie pairs. The data matrix is $Z\in\{1,\dots,5\}^{6040\times 3952}$ and mask matrix is $A\in\{0,1\}^{6040\times 3952}$. The data can be downloaded from https://grouplens.org/datasets/movielens/1m/. See https://files.grouplens.org/datasets/movielens/ml-1m-README.txt for the usage license.
D.4 Proposition 99
Data comes primarily from the Tax Burden on Tobacco compiled by Orzechowski and Walker [41] (ODC-By License). Using synthetic control methods, Abadie et al. construct a weighted combination of control states that closely resembles California’s pre-1988 characteristics and cigarette consumption patterns. The optimal weights produce a synthetic California primarily composed of Colorado (0.164), Connecticut (0.069), Montana (0.199), Nevada (0.234), and Utah (0.334), with all other states receiving zero weight. The treatment effect is estimated as the difference between actual California per-capita cigarette sales and those of synthetic California after Proposition 99’s implementation. By 2000, this analysis revealed that annual per-capita cigarette sales in California were approximately 26 packs lower than what they would have been without Proposition 99, representing about a 25% reduction in cigarette consumption. To validate these findings, the authors conducted placebo tests by applying the same methodology to states not implementing tobacco control programs, confirming that California’s reduction was unusually large and statistically significant (p = 0.026).
Proposition 99, the California Tobacco Tax and Health Protection Act of 1988, dataset spans from 1970 to 2000, providing 19 years of pre-intervention data before Proposition 99 was implemented in 1988 and 12 years of post-intervention data. It provides annual state-level cigarette consumption measured as per capita cigarette sales in packs based on tax revenue data. This data serves as a real data benchmark for many of the variants of synthetic controls [2]. We use the CDC dataset for the Nearest Neighbors methods and only use the target variable (i.e., cigarette consumption measured in packs per capita), and the dataset from SyntheticControlMethods library [^3] https://github.com/OscarEngelbrektson/SyntheticControlMethods/tree/master (Apache-2.0 License) for the SC baseline, since it relies on additional covariates.
D.5 PromptEval
MMLU is a multiple choice Q&A benchmark with $57$ tasks, with a total of near $14$K examples [^4] https://github.com/felipemaiapolo/prompteval (MIT License). $15$ different models, e.g., variants of Llama 3 [39], Mistral [31] and Gemma [50]. The examples are fed into the models with $100$ different varying prompting templates. The prompt templates are created by traversing between $3$ node modules, namely a separator, a space and an operator (see [43, Alg. 3, App. J] for details), from which $100$ unique prompt templates are created. The unitxt [12] preprocessing library is used to construct the dataset and evaluation is done by LM-Eval-Harness [26] library. The number of examples differ per task and each examples are evaluated on a model (verifiable, so assigned 00 or $1$ for correctness) by wrapping the examples with $100$ different prompt templates.